# Neuromorphic Principles for Efficient Large Language Models on Intel Loihi 2

### v1

@@ Abstract
Large language models (LLMs) deliver impressive performance but require large amounts of energy. In this work, we present a MatMul‐free LLM architecture adapted for Intel’s neuromorphic processor, Loihi 2. Our approach leverages Loihi2’s support for low-precision, event-driven computation and stateful processing. Our hardware-aware quantized model on GPU demonstrates that a 370M-parameter MatMul-free model can be quantized with no accuracy loss. Based on preliminary results, we report up to 3$\times$ higher throughput with 2$\times$ less energy, compared to transformer-based LLMs on an edge GPU, with significantly better scaling. Further hardware optimizations will improve increase throughput and decrease energy consumption. These results show the potential of neuromorphic hardware for efficient inference and pave the way for efficient reasoning models capable of generating complex, long-form text rapidly and cost-effectively.
@@ 1 Introduction
Loihi 2 is optimized for sequence-based processing, element-wise recurrence, low-precision arithmetic, and weight sparsity, all of which are features of the MatMul-free model. These benefits have been demonstrated on signal processing tasks (Orchard et al., 2021b; Shrestha et al., 2024) with orders of magnitude better latency and efficiency than state-of-the-art solutions. This paper presents a work-in-progress of adapting and deploying the MatMul-free language model from Zhu et al. (2024) to Intel Loihi 2, opening a pathway that bridges neuromorphic computing with state-of-the-art efficient LLMs. This required a microcode implementation to map the MMF model to Loihi 2’s architecture, along with a detailed ablation study to evaluate the optimal bit-precision for all operators in the MMF model. We are able to run the MMF model fully on-chip, using using fixed-point arithmetic to optimize for energy and latency.
@@ 3 Model Adaptation for Loihi 2
The Loihi 2 implementation represents the model as a network of neurons interconnected via synapses. Each neuron is implemented as a simple microcode program that is executed asynchronously on one of 120 neuro cores on each Loihi 2 chip, after which its output is transmitted to other neurons through synaptic connections. A global time step is maintained through barrier synchronization between all neuro cores. Since each neuron only maintains its own state, aggregate operations–such as computing the sum of squares over an activation vector–must be realized by dedicated neurons that receive inputs from all neurons within the corresponding layer. This is the case for the RMSNorm, where the sum of squares over all neurons in a given layer is calculated, see Equation 2. Figure A.5 (right) illustrates the computational graph that implements the RMSNorm operation on Loihi 2. Layer and operator fusion are well-established strategies to minimize redundant computations, thereby reducing latency and enhancing energy efficiency. In our implementation, we perform Loihi-specific layer fusion to consolidate operations into a streamlined computational graph, as depicted in Figure A.5 (left). We also derive the fusion of two subsequent RMSNorm layers into a single operator for further acceleration, see Appendix A.3.
@@ 4 Results
We implemented a single block of the MatMul-free language model on a single Loihi 2 chip. We parallelize the workload to all 120 available neuro cores on the chip. Verification of the model on Loihi 2 indicates close alignment with the quantized PyTorch simulation. In Table 2, we contrast the estimated performance of the 370M MatMul-free model on Loihi 2 against transformer-based baselines running on an NVIDIA Jetson Orin Nano. Efficiency and throughput metrics for Loihi 2 are estimates based on a preliminary implementation that is not fully optimized, see Appendix A.4.1 for details. We compare the MatMul-free LLM on Loihi 2 against two similarly-sized transformer-based LLMs available on HuggingFace. As a Llama series representative model, we use Alireo-400M (Montebovi, 2024), a 400M parameter transformer-based LLM with 24 layers and a context window of 8,192. We further use Qwen2.5-500M (Yang et al., 2024; Qwen Team, 2024), a 500M parameter transformer-based LLM with 24 layers and a context window of 32,768. Both models run in half-precision (FP16). We did not benchmark the MatMul-free LLM or the Transformer baseline from Zhu et al. (2024) because the Jetson Orin Nano does not support Triton.
Table 2: Throughput and energy efficiency for two transformer-based language models running on the NVIDIA Jetson Orin Nano compared to our MatMul-free LM running on Intel’s Loihi 2, across different sequence lengths for prefill and generation. The best-performing sequence length for each model and metric is underlined. Metrics for Loihi 2 are based on preliminary experiments and subject to further performance optimization, see Appendix A.4.1. Gen: autoregressive generation, Prefill: prefill mode. Llama representative model from Montebovi (2024).
@@
The MatMul-free LM on Loihi 2 was characterized on an Oheo Gulch single-chip Loihi 2 system (N3C1 silicon) running NxKernel v0.2.0 and NxCore v2.5.8 (accessible to Intel Neuromorphic Research Community members). Transformer LMs were characterized on NVIDIA Jetson Orin Nano 8GB using the MAXN power mode running Jetpack 6.2, TensorRT 10.3.0, CUDA 12.4. Energy values include CPU_GPU_CV, SOC, and IO components as reported by jtop 4.3.0. Performance results are based on testing as of Jan 2025 and may not reflect all publicly available security updates. Results may vary.
@@
During prefill, Loihi 2 shows at least 2x higher throughput with approximately 2x less energy per token. During auto-regressive generation, the advantage of Loihi 2 grows to having almost 3x higher throughput with approximately 2x less energy per token. Due to the MatMul-free model’s subquadratic architecture, this advantage is expected to grow for longer sequence lengths. We highlight that power and throughput of the MatMul-free model on Loihi 2 is constant across sequence length. This is due to the linear scaling of the recurrent token mixer and the fact that Loihi 2 has its parameters and hidden states stored locally inside neuro-cores thus requiring significantly less memory movement than GPUs. Further optimizations are underway to increase throughput and reduce power consumption of the MatMul-free language model on Loihi 2, see Appendix A.4.
Table 2 highlights how transformers on an edge GPU are consistently slower and less efficient during generation, and scale infavorably to longer sequence lengths. During prefill, throughput and efficiency is low for shorter sequences for transformers due to under-utilization of the GPU, then reach optimal throughput and efficiency between 4000-8000 sequence length, after which throughput and efficiency declines. Table 3 also compares the MatMul-free model on Loihi against the same model running on an H100 GPU. Loihi 2 delivers 3×3$\times$3 × higher throughput during generation with at least 14×14$\times$14 × less energy per token. During prefill, the H100 delivers higher throughput than Loihi 2, but outperforms in energy efficiency only for large sequence lengths.
@@ 5 Conclusion
We demonstrated that neuromorphic principles and co-design can be leveraged to build efficient LLMs. By merging a MatMul‐free architecture with Intel’s Loihi 2–leveraging state-based computation, extreme quantization, and operator fusion–we built a strong 370M-parameter model with significantly improved throughput and energy efficiency. Our experiments reveal that the inherent parallelism and low-power processing of Loihi 2 translate into substantial gains in throughput and energy efficiency. On the hardware side, our results demonstrates the potential of neuromorphic processors as platforms for scalable efficient inference, suggesting that future architectures can be co-designed with model innovations to further push performance limits. Our approach offers a promising pathway to enable adaptive language processing without the prohibitive energy costs associated with traditional LLMs. Given the rising importance of reasoning models that require extended chain-of-thought rollouts (DeepSeek-AI et al., 2025), efficient and high-throughput autoregressive generation is more critical than ever. Our design excels especially in this mode, paving the way for scalable foundation models capable of reasoning faster and more efficiently.
@@ A.1 Loihi 2 Hardware Architecture
The architectural features of Loihi 2 offer unique opportunities to compress and optimize deep learning models. Neural networks running on Loihi must be quantized to low-precision using fixed-point arithmetic–8 bits for synaptic weights [^4] This is not a hard limits, as one can implement an $8n$-bit synapse through $n$ separate 8-bit synpses that are added together with different fixed-point exponents. and up to 32 bits for messages [^5] Local states are not restricted in precision, and one may also transmit messages with more than 32 bits in an analogous way to what is described above for synaptic weights.. Unlike GPUs, Loihi 2 is optimized for computations local within neurons, a common focus of neuromorphic processors. First, it allows fast and efficient updates of neuronal states with recurrent dynamics with minimal data movement, due to the local memory of each neuro-core. Second, the asynchronous event-driven architecture of Loihi 2 allows it to efficiently process unstructured sparse weight matrices. Third, the neuro cores can leverage sparsified activation between neurons, as the asynchronous communication transfers only non-zero messages.
@@ A.1.1 Execution Modes on Loihi 2
Loihi 2’s asynchronous architecture allows to trade off between throughput and latency, as illustrated in Figure 3. For optimal throughput, new input is provided every time step and forwarded through the neuronal layers in a pipelined mode. For optimal latency we use fall-through mode, where new input is injected only once the previous input has been processed by, or fallen through, the network as fast as possible. Naturally, when processing a long sequence of prefill text in an LLM, we use pipeline mode for optimal throughput. When generating new text in auto-regressive generation of an LLM, we have to wait for the token at time $t$ to be output before we can begin processing the next token at time $t+1$, thus making this a natural fit for Loihi 2’s fall-through mode.
@@ A.4 Detailed Hardware Results
Table 3 shows extended results for the comparison of the MatMul-free LLM on Loihi 2 and transformer-based LLMs on the NVIDIA Jetson Orin Nano, also including results for the MatMul-free LLM on an H100 GPU and the Transformer++ baseline from Zhu et al. (2024) on an H100 GPU. The results demonstrate that the MatMul-free LLM on GPU improves in throughput and efficiency with longer sequence lengths, due to linear scaling of the token mixing and better utilization of the GPU at higher sequence lengths. In contrast, the transformer++ baseline on GPU increases in throughput for short sequence lengths due to better utilization of the hardware, and then deteriorates in throughput and efficiency for even longer sequences because of the quadratic scaling of self-attention.
Table 3: Throughput and energy efficiency for two transformer-based language models running on the NVIDIA Jetson Orin Nano compared to our MatMul-free LM running on Intel’s Loihi 2, across different sequence lengths for prefill and generation. The best-performing sequence length for each model and metric is underlined. Metrics for Loihi 2 are based on preliminary experiments and subject to further performance optimization, see Appendix A.4.1. Gen: autoregressive generation, Prefill: prefill mode. Llama representative model from Montebovi (2024).
Sequence length 500 1000 4000 8000 16000 500 1000 4000 8000 16000
Generate MMF (370M) Loihi 2 41.5 41.5 41.5 41.5 41.5 405 405 405 405 405
MMF (370M) H100 13.4 13.3 13.5 13.2 13.5 10.1k 10.1k 10.0k 9.9k 9.8k
TF++ (370M) H100 22.4 22.9 21.7 21.3 20.9 5.5k 5.6k 6.2k 6.8k 8.2k
Llama (400M) Jetson 14.3 14.9 14.7 15.2 12.8 723 719 853 812 1.2k
Qwen2 (500M) Jetson 13.4 14.0 14.1 15.4 12.6 791 785 912 839 1.2k
Prefill MMF (370M) Loihi 2 6632 6632 6632 6632 6632 3.7 3.7 3.7 3.7 3.7
MMF (370M) H100 11.4k 13.1k 30.6k 51.6k 84.6k 6.1 5.3 2.5 1.4 0.9
TF++ (370M) H100 21.6k 32.7k 44.3k 55.4k 60.5k 11.3 7.3 5.4 4.3 3.8
Llama (400M) Jetson 849 1620 3153 2258 1440 11.7 7.8 6.8 7.6 11.5
Qwen2 (500M) Jetson 627 909 2639 3861 3617 17.9 13.9 6.7 4.4 5.3
The MatMul-free LM on Loihi 2 was characterized on an Oheo Gulch single-chip Loihi 2 system (N3C1 silicon) running NxKernel v0.2.0 and NxCore v2.5.8 (accessible to Intel Neuromorphic Research Community members). Transformer LMs were characterized on NVIDIA Jetson Orin Nano 8GB using the MAXN power mode running Jetpack 6.2, TensorRT 10.3.0, CUDA 12.4. Energy values include CPU_GPU_CV, SOC, and IO components as reported by jtop 4.3.0. Performance results are based on testing as of Jan 2025 and may not reflect all publicly available security updates. Results may vary.
@@ A.4.1 Detailed Loihi 2 Results
The MatMul-free LM on Loihi 2 was characterized on (1) Oheo Gulch single-chip Loihi 2 system, (2) Alia Point 32-chip Loihi 2 system (both: N3C1 silicon) running NxKernel v0.2.0 and NxCore v2.5.8 (accessible to Intel Neuromorphic Research Community members). Transformer LMs were characterized on NVIDIA Jetson Orin Nano 8GB using the MAXN power mode running Jetpack 6.2, TensorRT 10.3.0, CUDA 12.4. Energy values include CPU_GPU_CV, SOC, and IO components as reported by jtop 4.3.0. Performance results are based on testing as of Jan 2025 and may not reflect all publicly available security updates. Results may vary.

### v2

@@ Abstract
Large language models (LLMs) deliver impressive performance but require large amounts of energy. In this work, we present a MatMul‐free LLM architecture adapted for Intel’s neuromorphic processor, Loihi 2. Our approach leverages Loihi 2’s support for low-precision, event-driven computation and stateful processing. Our hardware-aware quantized model on GPU demonstrates that a 370M-parameter MatMul-free model can be quantized with no accuracy loss. Based on preliminary results, we report up to 3$\times$ higher throughput with 2$\times$ less energy, compared to transformer-based LLMs on an edge GPU, with significantly better scaling. Further hardware optimizations will increase throughput and decrease energy consumption. These results show the potential of neuromorphic hardware for efficient inference and pave the way for efficient reasoning models capable of generating complex, long-form text rapidly and cost-effectively.
@@ 1 Introduction
Loihi 2 is optimized for sequence-based processing, element-wise recurrence, low-precision arithmetic, and weight sparsity, all of which are features of the MatMul-free model. These benefits have been demonstrated on signal processing tasks (Orchard et al., 2021b; Shrestha et al., 2024) with orders of magnitude better latency and efficiency than state-of-the-art solutions. This paper presents a work-in-progress of adapting and deploying the MatMul-free language model from Zhu et al. (2024) to Intel Loihi 2, opening a pathway that bridges neuromorphic computing with state-of-the-art efficient LLMs. This required a microcode implementation to map the MMF model to Loihi 2’s architecture, along with a detailed ablation study to evaluate the optimal bit-precision for all operators in the MMF model. We are able to run the MMF model fully on-chip, using fixed-point arithmetic to optimize for energy and latency.
@@ 3 Model Adaptation for Loihi 2
The Loihi 2 implementation represents the model as a network of neurons interconnected via synapses. Each neuron is implemented as a simple microcode program that is executed asynchronously on one of 120 neuro cores on each Loihi 2 chip, after which its output is transmitted to other neurons through synaptic connections. A global time step is maintained through barrier synchronization between all neuro cores. Since each neuron only maintains its own state, aggregate operations–such as computing the sum of squares over an activation vector–must be realized by dedicated neurons that receive inputs from all neurons within the corresponding layer. This is the case for the RMSNorm, where the sum of squares over all neurons in a given layer is calculated, see Equation 2. Figure A.5 (right) illustrates the computational graph that implements the RMSNorm operation on Loihi 2. Layer and operator fusion are well-established strategies to minimize redundant computations (Waeijen et al., 2021; Niu et al., 2021), thereby reducing latency and enhancing energy efficiency. In our implementation, we perform Loihi-specific layer fusion to consolidate operations into a streamlined computational graph, as depicted in Figure A.5 (left). We also derive the fusion of two subsequent RMSNorm layers into a single operator for further acceleration, see Appendix A.3.
@@ 4 Results
We implemented a single block of the MatMul-free language model on a single Loihi 2 chip. We parallelize the workload to all 120 available neuro cores on the chip. Verification of the model on Loihi 2 indicates close alignment with the quantized PyTorch simulation.
Note: All current comparisons are performed with FP16 baselines on non-Loihi hardware.
We contrast the estimated performance of the 370M MatMul-free model on Loihi 2 against transformer-based baselines running on an NVIDIA Jetson Orin Nano. We selected the Jetson Orin Nano (8GB) as our comparison platform because it represents a state-of-the-art edge AI device with 1024 CUDA cores, 32 Tensor cores, and a maximum power consumption of 15W, making it a relevant benchmark for energy-efficient AI applications. The Orin Nano is designed specifically for edge deployment scenarios similar to those targeted by neuromorphic hardware, enabling a fair comparison between platforms intended for similar operational environments. Efficiency and throughput metrics for Loihi 2 are estimates based on a preliminary implementation that is not fully optimized, see Appendix A.4.1 for details. We compare the MatMul-free LLM on Loihi 2 against two similarly-sized transformer-based LLMs available on HuggingFace. As a Llama series representative model, we use Alireo-400M (Montebovi, 2024), a 400M parameter transformer-based LLM with 24 layers and a context window of 8,192. It should be noted that the Alireo model was trained specifically on Italian text, so its performance is not included in Table 1 as it would not be representative of a competitive general-purpose transformer-based model at this parameter scale. We further use Qwen2.5-500M (Yang et al., 2024; Qwen Team, 2024), a 500M parameter transformer-based LLM with 24 layers and a context window of 32,768 whose performance is also included in Table 1. Both models run in half-precision (FP16). We did not benchmark the MatMul-free LLM or the Transformer baseline from Zhu et al. (2024) because the Jetson Orin Nano does not support Triton.
Table 2 shows results for the comparison of the MatMul-free LLM on Loihi 2 and transformer-based LLMs on the NVIDIA Jetson Orin Nano, also including results for the MatMul-free LLM on an H100 GPU and the Transformer++ baseline from Zhu et al. (2024) on a single H100 GPU. The results demonstrate that the MatMul-free LLM on GPU improves in throughput and efficiency with longer sequence lengths, due to linear scaling of the token mixing and better utilization of the GPU at higher sequence lengths. In contrast, the transformer++ baseline on GPU increases in throughput for short sequence lengths due to better utilization of the hardware, and then deteriorates in throughput and efficiency for even longer sequences because of the quadratic scaling of self-attention.
Table 2: Throughput and energy efficiency for two transformer-based language models running on the NVIDIA Jetson Orin Nano and H100 compared to our MatMul-free LM running on Intel’s Loihi 2, across different sequence lengths for prefill and generation. The best-performing sequence length for each model and metric is underlined. Metrics for Loihi 2 are based on preliminary experiments and subject to further performance optimization, see Appendix A.4.1. Gen: autoregressive generation, Prefill: prefill mode. Llama representative model from Montebovi (2024).
@@
The MatMul-free LM on Loihi 2 was characterized on a 32-chip Alia Point Loihi 2 system (N3C1 silicon) running NxKernel v0.2.0 and NxCore v2.5.8 (accessible to Intel Neuromorphic Research Community members). Appendix A.4.1 compares results for single-chip and multi-chip scaling. Transformer LMs were characterized on NVIDIA Jetson Orin Nano 8GB using the MAXN power mode running Jetpack 6.2, TensorRT 10.3.0, CUDA 12.4. Energy values include CPU_GPU_CV, SOC, and IO components as reported by jtop 4.3.0. Performance results are based on testing as of Jan 2025 and may not reflect all publicly available security updates. Results may vary.
@@
During prefill, Loihi 2 shows at least $\mathbf{2\times}$ higher throughput with approximately $\mathbf{2\times}$ less energy per token. During auto-regressive generation, the advantage of Loihi 2 grows to having almost $\mathbf{3\times}$ higher throughput with approximately $\mathbf{2\times}$ less energy per token. Due to the MatMul-free model’s subquadratic architecture, this advantage is expected to grow for longer sequence lengths. We highlight that power and throughput of the MatMul-free model on Loihi 2 is constant across sequence length. This is due to the linear scaling of the recurrent token mixer and the fact that Loihi 2 has its parameters and hidden states stored locally inside neuro-cores thus requiring significantly less memory movement than GPUs. Further optimizations are underway to increase throughput and reduce power consumption of the MatMul-free language model on Loihi 2, see Appendix A.4.
Table 2 highlights how transformers on an edge GPU are consistently slower and less efficient during generation, and scale unfavorably to longer sequence lengths. During prefill, transformers show low throughput and efficiency for shorter sequences due to under-utilization of the GPU, then reach optimal performance between 4000-8000 tokens, after which both throughput and efficiency decline. Table 2 also compares the MatMul-free model on Loihi against the same model running on an H100 GPU. Loihi 2 delivers $\mathbf{3\times}$ higher throughput during generation with at least $\mathbf{14\times}$ less energy per token. During prefill, the H100 delivers higher throughput than Loihi 2, but outperforms in energy efficiency only for large sequence lengths.
Latency characteristics are particularly important for interactive applications at the edge. Our experiments reveal that for batch size 1, which is typical in edge deployment scenarios, the MatMul-free model on Loihi 2 demonstrates significant latency advantages, with a $\times6$ lower time-to-first token on a 500-token input sequence (99ms for the MatMul-free model on the Loihi 2 vs. 659ms for the Llama-style model on the Jetson). This advantage increases with sequence length due to the linear scaling properties of our approach versus the quadratic complexity of transformer models. For real-time applications like voice assistants or mobile chatbots, this latency reduction directly translates to more responsive user experiences while maintaining significantly lower energy consumption.
While we benchmarked against the Jetson Orin Nano as a representative edge GPU platform, we expect our approach to show similar advantages compared to other edge computing solutions. Platforms such as mobile SoCs, FPGAs, and various edge TPUs all face similar challenges with transformer-based models: quadratic scaling with sequence length and significant memory movement costs. Since our neuromorphic approach addresses both fundamental limitations through linear scaling and closer compute-memory integration, we anticipate the relative throughput and energy efficiency advantages to persist across other edge computing architectures. The benefits would be most pronounced for resource-constrained IoT or mobile platforms where energy efficiency is paramount, with potentially more modest gains against specialized NPUs that have already optimized for recurrent operations.
@@ 5 Conclusion
We demonstrated that neuromorphic principles and co-design can be leveraged to build efficient LLMs. By merging a MatMul‐free architecture with Intel’s Loihi 2–leveraging state-based computation, extreme quantization, and operator fusion–we built a strong 370M-parameter model with significantly improved throughput and energy efficiency. Our experiments reveal that the inherent parallelism and low-power processing of Loihi 2 translate into substantial gains in throughput and energy efficiency.
The key innovations of our approach include: (1) the first demonstration of a modern LLM architecture on neuromorphic hardware, establishing a pathway for efficient AI at the edge; (2) a hardware-aware quantization methodology that maintains model accuracy while enabling fixed-point computation; (3) a novel microcode implementation of the MatMul-free architecture that exploits Loihi 2’s asynchronous, event-driven computing paradigm; and (4) custom operator fusion techniques specifically designed for neuromorphic computation, including our double RMSNorm derivation. Unlike previous approaches that targeted specific neural primitives for neuromorphic systems, our work shows that complete, competitive language models can be adapted to leverage the unique characteristics of neuromorphic hardware while maintaining performance.
On the hardware side, our results demonstrate the potential of neuromorphic processors as platforms for scalable efficient inference, suggesting that future architectures can be co-designed with model innovations to further push performance limits. Our approach offers a promising pathway to enable adaptive language processing without the prohibitive energy costs associated with traditional LLMs. Given the rising importance of reasoning models that require extended chain-of-thought rollouts (DeepSeek-AI et al., 2025), efficient and high-throughput autoregressive generation is more critical than ever. Our design excels especially in this mode, paving the way for scalable foundation models capable of reasoning faster and more efficiently.
@@ A.1 Loihi 2 Hardware Architecture
The architectural features of Loihi 2 offer unique opportunities to compress and optimize deep learning models. Neural networks running on Loihi must be quantized to low-precision using fixed-point arithmetic–8 bits for synaptic weights [^4] This is not a hard limit, as one can implement an $8n$-bit synapse through $n$ separate 8-bit synapses that are added together with different fixed-point exponents. and up to 32 bits for messages [^5] Local states are not restricted in precision, and one may also transmit messages with more than 32 bits in an analogous way to what is described above for synaptic weights.. Unlike GPUs, Loihi 2 is optimized for computations local within neurons, a common focus of neuromorphic processors. First, it allows fast and efficient updates of neuronal states with recurrent dynamics with minimal data movement, due to the local memory of each neuro-core. Second, the asynchronous event-driven architecture of Loihi 2 allows it to efficiently process unstructured sparse weight matrices. Third, the neuro cores can leverage sparsified activation between neurons, as the asynchronous communication transfers only non-zero messages.
@@ A.1.1 Execution Modes on Loihi 2
Loihi 2’s asynchronous architecture enables a trade off between throughput and latency, as illustrated in Figure 3. For optimal throughput, new input is provided every time step and forwarded through the neuronal layers in a pipelined mode. For optimal latency we use fall-through mode, where new input is injected only once the previous input has been processed by, or fallen through, the network as fast as possible. Naturally, when processing a long sequence of prefill text in an LLM, we use pipeline mode for optimal throughput. When generating new text in auto-regressive generation of an LLM, we have to wait for the token at time $t$ to be output before we can begin processing the next token at time $t+1$, thus making this a natural fit for Loihi 2’s fall-through mode.
@@ A.4.1 Detailed Loihi 2 Results
The MatMul-free LM on Loihi 2 was characterized on (1) Oheo Gulch single-chip Loihi 2 system, (2) Alia Point 32-chip Loihi 2 system (both: N3C1 silicon) running NxKernel v0.2.0 and NxCore v2.5.8 (accessible to Intel Neuromorphic Research Community members). Detailed results for single-chip and multi-chip scaling are presented in Appendix A.4.1. Transformer LMs were characterized on NVIDIA Jetson Orin Nano 8GB using the MAXN power mode running Jetpack 6.2, TensorRT 10.3.0, CUDA 12.4. Energy values include CPU_GPU_CV, SOC, and IO components as reported by jtop 4.3.0. Performance results are based on testing as of Jan 2025 and may not reflect all publicly available security updates. Results may vary.